# A Survey on Self-play Methods in Reinforcement Learning

### v1

@@ IV-B2 DouDiZhu
DouDizhu (a.k.a. Fight the Landlord) is a strategic Chinese card game for three players. In this game, one player takes on the role of the landlord and competes against the other two players, who are the peasants. These two roles have distinct objectives: the landlord’s goal is to outplay the peasants, whereas the peasants work together to defeat the landlord. The essence of the game lies in its blend of strategy, cooperation among peasants, and competition between the two factions. The game is played in two main stages: the bidding stage and the card play stage. During the bidding stage, players vie to become the landlord through a process that involves assessing the strength of their cards. The outcome of this stage determines the game’s dynamics, with the landlord playing against the peasants. The card play stage is where the main action unfolds. Players take turns playing cards in various combinations, intending to be the first to empty their hands. The game’s complexity arises from its requirement for strategic thinking, both in playing cards effectively and in the peasants’ need to cooperate seamlessly. DouDiZhu is characterized by imperfect information, as players can only see their own cards and the cards played, adding a layer of unpredictability and strategic depth to the game. Moreover, with an estimated $10^{76}\sim 10^{106}$ possible game states [117] and an action space comprising $27472$ possible moves [118], DouDiZhu presents a vast landscape of strategic possibilities.
DeltaDou [119], built upon the AlphaZero framework [10], is the first algorithm to achieve expert-level performance in DouDizhu although it only considers the card play stage. It enhances the original framework by incorporating asymmetric MCTS to improve exploration in games characterized by imperfect information. Additionally, DeltaDou employs policy-value networks specially designed with input and output encodings to accelerate the training process. A notable feature of DeltaDou is its ability to infer the cards of other players using Bayesian methods with the help of the policy-value networks. DouZero [120] reduces training costs by opting for a sampling method over the tree search approach. It utilizes Monte Carlo methods combined with deep neural networks. More specifically, it works similarly to independent RL self-play to generate data, which is then stored in buffers. Subsequently, data is sampled from the buffers to update the neural networks. Remarkably, DouZero surpasses DeltaDou in performance, demonstrating the effectiveness of its methodology in mastering. Additionally, DouZero employs supervised learning to train a bidding network used during the bidding stage, utilizing expert data for this purpose. Based on DouZero, DouZero+ [121] enhances the original algorithm by incorporating opponent modeling, which assists the agent in making more informed decisions. Additionally, DouZero+ introduces coach-guided learning, aiming to ensure a balanced intensity of initial hand cards among the three players to further accelerate the training process. PerfectDou [118] leverages perfect information distillation within a Perfect-Training-Imperfect-Execution (PTIE) framework. In this approach, the value network is trained using both perfect and imperfect features, while the policy network is trained solely on imperfect features. The networks are refined through independent RL self-play, utilizing Proximal Policy Optimization (PPO) [122] with Generalized Advantage Estimation (GAE) [123]. Notably, PerfectDou does not incorporate the bidding stage into its model, thereby eliminating the need for expert data in this phase. PerfectDou achieves better performance than DouZero.
IV-B3 Mahjong
Mahjong, renowned for its complex interplay of skill, strategy, and chance, has evolved into various global variants, including the popular Japanese version known as Riichi Mahjong. This game is typically played by four participants who must navigate both the visible aspects of the game, such as discarded tiles, and the hidden elements, like their own hand and the unseen public tiles. The strategic depth and complexity of Mahjong pose significant challenges for artificial intelligence. Despite ongoing research, many AI implementations have yet to reach expert human levels [124, 125, 126]. AI systems must make decisions based on incomplete information and adapt strategies dynamically in response to the actions of multiple opponents, which complicates the decision-making process. Additionally, the game’s intricate winning rules and the enormous number of possible game states, estimated at around $10^{169}$ [117], add layers of complexity. These factors make Mahjong a particularly challenging domain for AI development, requiring sophisticated approaches to handle its unique demands effectively.
Suphx [127] is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, specifically 10 dan on Tenhou [128], the most popular online Mahjong platform. Initially, Suphx employs supervised learning, utilizing expert data to train its model. It then advances its capabilities through self-play RL, employing several innovative techniques to enhance its effectiveness. These techniques include global reward prediction, which provides a continuous reward throughout the game, and oracle guiding, which offers perfect information at the start of self-play to guide strategy development. Additionally, parametric Monte Carlo policy adaptation is used to refine the policy further. Similarly, NAGA [129], developed by Dwango Media Village, and LuckyJ [130], designed by Tencent, have also achieved the rank of 10 dan on Tenhou. Furthermore, LuckyJ has even defeated human professional players,
IV-C Video Games
@@
In contrast to traditional board games and card games, video games often feature real-time actions, long time horizons and a higher level of complexity due to the broader range of possible actions and observations. Here, we illustrate some representative video games to demonstrate how self-play has advanced AI within these games.
IV-C1 StarCraft II
StarCraft is a real-time strategy (RTS) game developed and published by Blizzard Entertainment. It showcases three distinct species: the Terrans, Zerg, and Protoss, each with unique units and strategic options that enhance the gameplay experience. Renowned for its balanced gameplay, strategic depth, and challenging multiplayer features, the game tasks players with gathering resources, constructing bases, and building armies. Victory requires meticulous planning and tactical execution, with defeat occurring when a player loses all their buildings.
@@ IV-C1 StarCraft II
AlphaStar [15], a significant advancement in AI, dominates the 1v1 mode competitions in StarCraft II and has defeated professional players. Its framework bears similarities to that of AlphaGo [8], utilizing supervised learning initially to train the policy with data from human experts. Subsequently, it uses end-to-end RL and a hierarchical self-play method to further train the networks. More specifically, it divides all the agents into three types: main agents, league exploiters, and main exploiters. Additionally, it maintains a policy pool of past players that records all these types of agents. Main agents engage in both FSP and PFSP, competing against main agents themselves and other agents in the policy pool. They are periodically added to the pool and never reset. League exploiters use PFSP to play against all policy pool agents, added to the pool if they show a high win rate and potentially reset to expose global blind spots. Main exploiters only compete with main agents to improve their robustness and are added to the pool after achieving a high win rate or certain training steps, and are reset upon each addition. Most importantly, among those three agent types, the main agent is the core agent and embodies the final AlphaStar strategy. However, the training computation for AlphaStar is extensive. Further studies [131, 132, 133] have enhanced the league self-play training procedure by introducing innovative techniques to reduce computations.
IV-C2 MOBA Games
@@
Multiplayer Online Battle Arena games (MOBA) are a popular genre of video games that blend RTS with role-playing elements. In typical MOBA games, two teams of players control their own unique characters, known as heroes and compete to destroy the opposing team’s main structure, often called the base. Each hero has distinct abilities and plays a specific role within the team, such as Warrior, Tank or Support. Managing multiple lanes and battling under the fog of war, which obscures parts of the map are key aspects of gameplay. Popular examples of MOBA games include League of Legends, Dota 2 and Honor of Kings. These games are known for their complex strategic depth, decision-making under conditions of imperfect information and emphasis on teamwork and player skill.
@@
OpenAI Five [14] defeated the world champion team in a simplified version of Dota 2 that featured a limited pool of heroes and certain banned items. It employs distributed RL using PPO [122] along with GAE [123] to scale up training. Each player on a team shares the same policy network and receives shared observation inputs, with the only difference being a unique identity input for each player. The training process utilizes self-play, with an 80% chance of engaging in vanilla self-play and a 20% chance of using a self-play technique similar to PFSP used in AlphaStar [15]. Here, the probabilities of selecting each policy from the policy pool vary according to its quality score, which is continually updated based on competition results throughout the training. Higher quality scores increase the likelihood of a policy being selected. A new agent is added to the pool every 10 iterations. Similarly to AlphaStar, OpenAI Five requires extensive training resources. Thus, they developed surgery tools to resume training with minimal performance loss to manage continual changes in the environment and codes during development. Additionally, during training, OpenAI Five employs TrueSkill [40] to evaluate agent performance.
@@
Another notable MOBA game, especially popular in China, is Honor of Kings. The 1v1 mode is conquered by [134], which boasts a significant winning rate against top professional players. This success can be attributed to well-designed systems tailored for large-scale RL training and sophisticated algorithm designs that handle complex control challenges effectively. A key element of its RL training methodology is the use of $\delta$-uniform Self-play. The 5v5 mode was later mastered by [135] as well. Unlike OpenAI Five, this work expands the hero pool to 40 heroes, which exponentially increases the possible combinations of lineups and poses a learning collapse problem in RL training. It proposes curriculum self-play learning (CSPL) to mitigate this issue. Specifically, the training process is divided into three stages, drawing on insights from curriculum learning. The first stage involves training fixed lineups through self-play and utilizing human data to keep the two teams balanced to aid policy improvements. The second stage employs multi-teacher policy distillation to produce a distilled model. The final stage uses this distilled model as the initial model for self-play with randomly picked lineups. This approach defeats professional player teams. Additionally, the self-play generated data is used to learn effective lineup drafting by utilizing MCTS and neural networks [136].
IV-C3 Google Research Football
@@
Google Research Football (GRF) [137] is an open-source football simulator that emphasizes high-level actions. It initially offers two scenarios: the football benchmark and the football academy with 11 specific tasks. Here, we focus exclusively on the football benchmark because it presents a more complex scenario that better demonstrates the effects of self-play, as opposed to testing on specially designed tasks. GRF is particularly challenging due to the need for cooperation among teammates and competition against opposing teams. Additionally, it features long trajectories with 3000 steps per round, stochastic transitions, and sparse rewards, all of which contribute to its complexity.
@@
WeKick [138] claimed victory in the GRF competition on Kaggle [139], which modified the game dynamics by allowing competitors to control just one player, either the ball carrier on offense or the nearest defender on defense. WeKick employed self-play strategies similar to those used in league training [15]. It initializes its opponent policy pool using strategies developed through reward shaping via RL, and through Generative Adversarial Imitation Learning [140], which facilitated learning from the tactics of other teams.
@@
Further research delves into the full football game rather than controlling only one player in the team. Team-PSRO [141] extends PSRO [42] to team games, demonstrating approximate TMECor [34] and outperforming self-play in the 4v4 version of the full GRF. In the context of the 11v11 version of the full GRF, where the goalkeeper is rule-based controlled, TiKick [142] employs imitation learning to glean insights from data gathered by WeKick’s self-play, resulting in a multi-agent AI through distributed offline RL. Moreover, Fictitious Cross-Play (FXP) [143] proposes two populations: the main population and the counter population. Policies in the counter population improve solely by cross-playing with policies in the main population as opponents, while policies in the main population engage in playing with policies from both populations. FXP achieves a win rate of over 94% against TiKick in the 11v11 version of the full GRF. TiZero [144], a follow-up to TiKick, combines curriculum learning with FSP and PFSP [15] to avoid expert data reliance, achieving a higher TrueSkill rating [40] than TiKick.
@@
Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero [10] Independent RL $\checkmark$ $\times$
MuZero [11] Independent RL $\checkmark$ $\times$
@@
AlphaGo Zero [9] Independent RL $\checkmark$ $\times$
AlphaZero [10] Independent RL $\checkmark$ $\times$
MuZero [11] Independent RL $\checkmark$ $\times$
@@
DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou [119] Independent RL $\checkmark$ $\times$
DouZero [120] Independent RL $\times$ $\checkmark$
PerfectDou [118] Independent RL $\times$ $\times$
Mahjong 4 $\times$ $10^{169}$ Suphx [127] Independent RL $\checkmark$ $\checkmark$
@@
Dota 2 10 $\times$ OpenAI Five [14] Naive SP & PFSP $\times$ $\times$
@@
Table III: Overview of Representative Studies in Empirical Analysis.
V Open Problems and Future Work
Self-play approaches have demonstrated superior performance due to their unique iterative learning processes and ability to adapt to complex environments. However, there are still several areas that require further research and development.
V-A Theoretical Foundation
Although NE has been shown to exist in games with finite players and finite actions [145], computing NE in larger games remains challenging and consequently, many studies aim to achieve approximate NE [146]. However, in some cases, even computing an approximate NE is difficult [65]. Some research has resorted to higher levels of equilibrium, such as CE [47] and $\alpha$-rank [46].
@@ V-A Theoretical Foundation
Although many algorithms are developed with theoretical foundations in game theory, there is often a gap when it comes to applying these algorithms to complex real-world scenarios. For example, although AlphaGo [8], AlphaStar [147] and OpenAI Five [14] have achieved empirical success, they lack formal game theory proofs behind their effectiveness. Future work should focus on bridging this gap by combining empirical success with theoretical validation. This could involve developing new algorithms that are both theoretically sound and practically effective in complex scenarios or proving the theoretical underpinnings of existing successful algorithms in complex environments.
V-B Non-stationarity of the Environment
@@
The strategies of the opponent players evolve as training progresses, and the opponents are a vital factor of the environment in the self-play framework. This evolution can cause the same strategy to lead to different results over time, creating a non-stationary environment. This problem is also shared by the MARL area. Future research should aim to develop algorithms that are more robust and can adapt to changing conditions. For example, incorporating opponent modeling into self-play [121] can help agents anticipate changes in opponent strategies and adjust their own strategies proactively, making them more robust to environmental changes.
V-C Scalability and Training Efficiency
The scalability of self-play methods faces significant challenges as the number of teams and players within those teams increases. As the number of participants grows, the complexity of interactions explodes. For example, in OpenAI Five [14], the hero pool size is limited to only 17 heroes. MOBA AI [135] extends this to a 40-hero pool with the help of curriculum learning, but it still cannot cover the entire hero pool available in the actual game. One potential solution is to leverage the inherent connections among players to optimize the learning process. For instance, using graph-based models to represent and exploit the relationships between players can help manage and reduce the complexity of large-scale multi-agent environments.
@@ V-C Scalability and Training Efficiency
These scalability issues are fundamentally rooted in the limited training efficiency of self-play methods, which involve both computational and storage aspects. Computational efficiency is a limiting factor due to the iterative nature of self-play, where agents repeatedly play against themselves or past versions, requiring extensive computational resources. Moreover, although forming more complex populations and competitive mechanisms [147] can enhance the intensity and quality of training, it further exacerbates the demand for computational resources. Techniques such as parallel computing, distributed learning, and more efficient neural network architectures could be explored to address these challenges. Additionally, self-play is storage-intensive, as it requires maintaining a policy pool. Even when using a shared network architecture, storing the parameters of large models can be problematic. This issue is particularly pronounced in regret-minimization-based self-play algorithms, which need to store the regrets for each information set and potential action. Managing both the computational load and storage requirements is essential for improving the overall training efficiency and scalability of self-play methods.
V-D With Large Language Models

### v2

@@ 4.2.2 DouDiZhu
DouDizhu (a.k.a. Fight the Landlord) is a three-player strategic card game. In this game, one player takes on the role of the landlord and competes against the other two players, the peasants. The game is played in two main stages: the bidding stage and the card play stage. During the bidding stage, players vie to become the landlord. During the card play stage, players take turns playing cards in various combinations, intending to be the first to empty their hands. DouDiZhu is characterized by imperfect information, as players can only see their own cards and the cards played. The essence of the game lies in cooperation among peasants and competition between the two factions. It has an estimated $10^{76}\sim 10^{106}$ possible game states [147] and an action space comprising $27472$ possible moves [141].
DeltaDou [61] is the first algorithm to achieve expert-level performance in DouDizhu. Similar to AlphaZero [116], DeltaDou utilizes vanilla self-play to generate samples to train policy and value networks. DouZero [148] also uses vanilla self-play to generate data. It reduces training costs by opting for a sampling method over the tree search approach. Based on DouZero, DouZero+ [150] incorporates opponent modeling, assisting the agent in making more informed decisions. PerfectDou [141] utilizes vanilla self-play to train a Perfect-Training-Imperfect-Execution framework.
4.2.3 Mahjong
Mahjong has evolved into various global variants, including the famous Japanese version known as Riichi Mahjong. This game is typically played by four players who must navigate both the visible aspects of the game, such as discarded tiles, and the hidden elements, like their own hand and the unseen public tiles. The strategic depth and complexity of Mahjong pose significant challenges. Despite ongoing research, these findings have yet to reach expert human levels [71, 40, 93]. Difficulties are navigating incomplete information, dynamically adapting strategies to multiple opponents, and contending with complex winning rules and an enormous number of possible game states, estimated at $10^{169}$ [147]. Suphx [77] is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, precisely 10 dan on Tenhou [127], the most popular online Mahjong platform. Initially, Suphx employs supervised learning, utilizing expert data to train its model. It then advances its capabilities through vanilla self-play combined with RL. Similarly, NAGA [128], developed by Dwango Media Village, and LuckyJ [125], designed by Tencent, have also achieved the rank of 10 dan on Tenhou. Furthermore, LuckyJ has even defeated human professional players.
4.3 Video Games
@@
In contrast to traditional board games and card games, video games often feature real-time actions, long trajectories, and increased complexity stemming from a broader range of actions and observations. Here, we illustrate some representative video games to demonstrate how self-play has advanced AI within these games.
4.3.1 StarCraft II
@@
StarCraft is a real-time strategy (RTS) game. It has three distinct species: the Terrans, Zerg, and Protoss, each with unique units and strategic options that enhance the gameplay complexity. Renowned for its balanced gameplay, strategic depth, and intense competitiveness, the game challenges players to gather resources, construct bases, and build armies. Victory requires meticulous planning and tactical execution, with defeat occurring when a player loses all their buildings. AlphaStar [129] dominates the 1v1 mode competitions in StarCraft II and has defeated professional players. Its framework is similar to AlphaGo [115], initially utilizing supervised learning to train the policy with expert data. Subsequently, it uses a hierarchical self-play method combined with end-to-end RL to train the networks. More specifically, The proposed self-play method divides all the agents into three types: main agents, league exploiters, and main exploiters. It maintains a policy pool of past players that records all these types of agents. Main agents engage in traditional self-play algorithms like FSP and PFSP, competing against main agents themselves and other agents in the policy pool. They are periodically added to the pool and never reset. League exploiters also use a traditional self-play method PFSP to play against all policy pool agents, added to the pool if they show a high win rate and potentially reset to expose global blind spots. Main exploiters only compete with main agents to improve their robustness, are added to the pool after achieving a high win rate or specific training steps, and are reset upon each addition. Among those three agent types, the main agent is the core agent and embodies the final AlphaStar strategy. However, the introduction of three agent types significantly increases the training computation. Further studies [48, 131, 57] have enhanced the league self-play training procedure.
4.3.2 MOBA Games
@@
Multiplayer Online Battle Arena (MOBA) games are a popular video game genre that blends RTS with role-playing elements. In typical MOBA games, two teams of players control their unique characters, known as heroes, and compete to destroy the opposing team’s main structure, often referred to as the base. Each hero has distinct abilities and plays a specific role within the team, such as Warrior, Tank, or Support. Managing multiple lanes and battling under the fog of war, which obscures parts of the map, are critical aspects of the gameplay. Popular examples of MOBA games include League of Legends, Dota 2, and Honor of Kings.
@@
OpenAI Five [7] defeated the world champion team in a simplified version of Dota 2 that featured a limited pool of heroes and certain banned items. The training process introduces a self-play method that combines two traditional self-play algorithms: with an 80% probability of engaging in vanilla self-play and a 20% probability of employing a technique similar to the PFSP used in AlphaStar [129]. This technique selects each policy from the policy pool based on its quality score, which is continuously updated from competition results throughout training. Higher quality scores increase the likelihood of a policy being selected. OpenAI Five also requires extensive training resources.
@@
Another notable MOBA game, especially viral in China, is Honor of Kings. The 1v1 mode is conquered by [144], which boasts a significant winning rate against top professional players. It utilizes a traditional self-play method $\delta$-uniform self-play to train the RL networks. The 5v5 mode was later mastered by [143] as well. Unlike OpenAI Five, this work expands the hero pool to 40 heroes, exponentially increasing the possible combinations of lineups. It proposes a new self-play method referred to as curriculum self-play learning (CSPL). Specifically, the training process is divided into three stages. The first stage involves training fixed lineups through self-play and utilizing human data to balance the two teams to aid policy improvements. The second stage employs multi-teacher policy distillation to produce a distilled model. The final stage uses this distilled model as the initial model for another round of self-play with randomly picked lineups. This approach defeats professional player teams. The self-play generated data is also used to learn compelling lineup drafting by utilizing MCTS and neural networks [25].
4.3.3 Google Research Football
@@
Google Research Football (GRF) [70] is an open-source football simulator emphasizing high-level actions. It initially offers two scenarios: the football benchmark and the football academy with 11 specific tasks. Here, we focus exclusively on the football benchmark because it presents a more complex scenario that better demonstrates the effects of self-play. GRF is particularly challenging due to the need for cooperation among teammates and competition against opposing teams. It features long trajectories with 3000 steps per round, stochastic transitions, and sparse rewards.
@@
WeKick [153] claimed victory in the GRF competition on Kaggle [45], which simplifies the game dynamics by allowing competitors to control only one player, either the ball carrier on offense or the nearest defender on defense. WeKick employed self-play strategies similar to those used in league training [129]. It initializes its opponent policy pool using strategies developed through RL and Generative Adversarial Imitation Learning (GAIL) [55] to facilitate the training process.
@@
Further research delves into the full football game rather than the simplified version. Team-PSRO [91] extends PSRO series of self-play algorithms to team games, outperforming baselines in the 4v4 version of the full GRF. In the context of the 11v11 version, where the goalkeeper is rule-based controlled, TiKick [58] utilizes vanilla self-play to collect samples and then employs imitation learning. Fictitious Cross-Play (FXP) [137] proposes a new self-play method similar to AlphaStar [129]. It introduces two populations: the main population and the counter population. Policies in the counter population improve solely by cross-playing with policies in the main population as opponents, while policies in the main population engage in playing with policies from both populations. FXP achieves a win rate of over 94% against TiKick. TiZero [78], a follow-up to TiKick, combines curriculum learning with FSP and PFSP [129] to avoid expert data reliance, achieving a higher TrueSkill rating [53] than TiKick.
Table 2: Overview of Representative Studies in Empirical Analysis.
@@
Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero [116] Vanilla SP $\checkmark$ $\times$
MuZero [112] Vanilla SP $\checkmark$ $\times$
@@
AlphaGo Zero [117] Vanilla SP $\checkmark$ $\times$
AlphaZero [116] Vanilla SP $\checkmark$ $\times$
MuZero [112] Vanilla SP $\checkmark$ $\times$
@@
DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou [61] Vanilla SP $\checkmark$ $\times$
DouZero [148] Vanilla SP $\times$ $\checkmark$
PerfectDou [141] Vanilla SP $\times$ $\times$
Mahjong 4 $\times$ $10^{169}$ Suphx [77] Vanilla SP $\checkmark$ $\checkmark$
@@
Dota 2 10 $\times$ OpenAI Five [7] Vanilla SP & PFSP $\times$ $\times$
@@
5 Open Problems and Future Work
Self-play methods have shown remarkable performance by leveraging iterative learning and adapting to complex environments. However, significant challenges and open questions remain, presenting opportunities for further research and development.
5.1 Theoretical Foundation
@@
Although NE has been shown to exist in games with finite players and finite actions [100], computing NE with self-play algorithms in larger games remains challenging and consequently, many studies aim to achieve approximate NE [76]. However, in some cases, even computing an approximate NE is difficult [31]. Some research has resorted to higher levels of equilibrium, such as CE [86] and $\alpha$-rank [96]. Although many algorithms are developed with theoretical foundations, there is often a gap in complex real-world scenarios. For example, although self-play algorithms like AlphaGo [115], AlphaStar [109], and OpenAI Five [7] achieve empirical success, they lack formal game theory proofs behind their effectiveness. Future work should focus on bridging this gap by developing new theoretically sound and practically effective algorithms or proving the theoretical underpinnings of existing successful self-play algorithms in complex environments.
5.2 Non-stationarity of the Environment
@@
In the self-play framework, the opponents are a vital component of the environment, and the strategies of the opponent players evolve as training progresses. This evolution can cause the same strategy to lead to different results over time, creating a non-stationary environment. This problem is also shared by the MARL area. Future research should aim to develop self-play algorithms that are more robust and can adapt to changing conditions. For example, incorporating opponent modeling [150] into self-play can help agents anticipate changes in opponent strategies and adjust their own strategies proactively, making them more robust to environmental changes.
5.3 Scalability and Training Efficiency
@@
The scalability of self-play methods faces significant challenges as the number of teams and players within those teams increases. As the number of participants grows, the complexity of interactions explodes. For example, in OpenAI Five [7], the hero pool size is limited to only 17 heroes. MOBA AI [143] extends this to a 40-hero pool with the help of curriculum learning, but it still cannot cover the entire hero pool available in the actual game. One potential solution is leveraging players’ inherent connections to optimize the learning process. For instance, using graph-based models to represent and exploit the relationships between players can help manage and reduce the complexity of large-scale multi-agent environments. These scalability issues are fundamentally rooted in the limited training efficiency of self-play methods from two aspects including computation and storage. The first issue is computational efficiency, which is induced by the iterative nature of self-play where agents repeatedly play against themselves or past versions. Moreover, although forming more complex populations and competitive mechanisms [109] can enhance the intensity and quality of training, it further exacerbates the demand for computational resources. Techniques such as parallel computing, distributed learning, and more efficient neural network architectures could be explored to address these challenges. The second issue is storage because self-play requires maintaining a policy pool. Even when using a shared network architecture, storing the parameters of large models can be problematic. This issue is particularly pronounced in regret-minimization-based self-play algorithms, which must store the regrets for each information set and potential action. Managing both the computational load and storage requirements is essential for improving the overall training efficiency and scalability of self-play methods.
5.4 With Large Language Models